The Cobb- Douglas production function is given by the form Y = ALαKβ

Where, Y = Output (mt/day), L = Labour (hours/mt) K = Capita (Rs/mt). By taking the natural logarithm of each term in the function, that is LnY = LnA+αLnL+βLnK. Using a standard multiple linear regression model the parameters of the function were obtained as follows;

Ln Y = 1.389 +0.8LnL + 0.6LnK

            (0.53)    (0.02)      (0.04)


1
Expert's answer
2021-09-01T11:51:41-0400

Elasticity of labour "=" "\frac{dY}{DL}" "\times" "\frac{L}{Y}"

Now;

In Y "=" 1.388+0.8lnL + 0.6ln K

Differentiating both sides w.r.t. labour we get:

"\frac{1}{Y}" "\times" "\frac{dY}{DL}=" "\frac{0.8}{L}"

"\frac{dY}{dL}=" "\frac{0.8Y}{L}"

Putting this value in elasticity equation:

E(L)="\frac{0.8Y}{L}" "\times" "\frac{L}{Y}"

E(L)=0.8

Hence the change in the rate of labor leads to 0.8 change in the level of output

Similarly,

Elasticity of labor ="\frac{dY}{dK}\times\frac{K}{Y}"

Differentiating output equation w.r.t to capital we get

"\frac{1}{Y}\times\frac{dY}{dK}=\frac{0.6}{K}"

Putting in elasticity equation we get;

E(K)="\frac{0.6Y}{K}\times\frac{K}{Y}"

E(K)=0.6

Hence the change in the rate of capital leads to 0.6 change in the level of output.